You write one Chih, Inc. April 120 put contract (equaling 100 shares) for a premium of $4. You hold the option until the expiration date, when Chih stock sells for $118 per share. You will realize a ______ on the investment.

Multiple Choice

  • $600 loss

    Incorrect
  • $200 profit

  • $200 loss

  • $300 profit
Answer #1

You are a writer of a put option. So premium is your income

Your income = 4 x 100 = 400

Now, you have negative payoff if price decreases and zero payoff if price increases beyond strike price 120.

Now, here price on expiry is 118, so payoff will be negative

Payoff = price on expiry - strike price = 118 - 120 = -2 x 100 = -200

Profit or loss = Payoff + premium = - 200 + 400 = 200

So there is a profit of 200

Answer : $200 profit (Thumbs up please)